# Using AI-Based Coding Assistants in Practice: State of Affairs, Perceptions, and Ways Forward

### v1

@@ Abstract
The last several years saw the emergence of AI assistants for code — multi-purpose AI-based helpers in software engineering. Their quick development makes it necessary to better understand how specifically developers are using them, why they are not using them in certain parts of their development workflow, and what needs to be improved.
In this work, we carried out a large-scale survey aimed at how AI assistants are used, focusing on specific software development activities and stages. We collected opinions of 481 programmers on five broad activities: (a) implementing new features, (b) writing tests, (c) bug triaging, (d) refactoring, and (e) writing natural-language artifacts, as well as their individual stages.
Our results show that usage of AI assistants varies depending on activity and stage. For instance, developers find writing tests and natural-language artifacts to be the least enjoyable activities and want to delegate them the most, currently using AI assistants to generate tests and test data, as well as generating comments and docstrings most of all. This can be a good focus for features aimed to help developers right now. As for why developers do not use assistants, in addition to general things like trust and company policies, there are fixable issues that can serve as a guide for further research, e.g., the lack of project-size context, and lack of awareness about assistants. We believe that our comprehensive and specific results are especially needed now to steer active research toward where users actually need AI assistants.
I Introduction
11footnotetext: The first two authors contributed equally to this work.
@@ I Introduction
Given their rapid growth, understanding software developers’ perceptions and needs regarding AI assistants is paramount in shaping the future of software development. Recent studies have started looking into why AI assistants are used, what their high-level drawbacks are [10, 11, 12], etc. However, the existing studies do not consider individual activities within the software development life-cycle and their specificities. Such comprehensive yet granular analysis will help us prioritize the research in the field by finding areas where AI assistants are already employed to focus on them in the short term and identify shortcomings to address them in the future. To the best of our knowledge, no study has yet been conducted on such an analysis.
@@ 3. RQ3
In terms of the general usage of AI assistants and their perception, our results show that 84.2% of the respondents occasionally or regularly use at least one of the tools in the given list, the most popular ones being ChatGPT (72.1%) and GitHub Copilot (37.9%). Also, the most positively rated quality of the code provided by AI assistants is usable (56% of respondents agreeing), while the most negatively rated is secure (only 23% agreeing).
Studying individual activities and stages demonstrated their differences. We found that 87.3% of respondents use AI assistants at least in some stages of Implementing new features, about three quarters use them at least in some stages of Writing tests, Refactoring, and Writing natural language artifacts, and only 65.5% use them for Bug triaging. When comparing individual stages of these activities, we found that developers use AI more at stages where generating or summarizing code is necessary. Developers use AI less for finding places in the code to apply the change or directly applying it to the code. Given the fact that writing tests and writing natural language artifacts were described as less enjoyable and more likely to be delegated, this represents areas which the researchers can focus on right now to bring direct value to the users.
Our thematic analysis of answers to the open-ended questions about the reasons for not using AI assistants revealed 20 distinct themes. The most popular ones are Lack of need for AI assistance, AI-generated output being inaccurate, User’s lack of trust and the desire to feel in control, and Lack of understanding context by AI assistant. While they are prevalent across all activities, some are more noticeable in specific ones, e.g., Lack of understanding context is important in Bug triaging, with respondents highlighting that finding bugs usually requires a deeper understanding of both code and business logic. We formulate a list of the most prevalent issues that can be addressed by the community in future work. The supplementary materials for the paper are available online and can be found on Zenodo [13].
@@
  • Usage patterns. We report how AI assistants are employed in the software development life-cycle of developers and which activities and individual stages are more prevalent in usage.
  • Areas of focus. Taking into account which activities the developers find less enjoyable and want to delegate, as well as what stages AI assistants are used on, we highlight areas where the research needs to focus on to bring value to users right now.
  • Reasons for not using AI. We provide 20 distinct themes that represent different reasons why developers are not using AI assistants and report the main ones for all activities.
  • Future work. Based on the prevalence of different reasons, we formulate the main areas of future work that are needed to overcome the shortcomings that the respondents describe.
The remainder of this paper is organized as follows. In Section II, we discuss the related work. In Section III, we describe how we developed and piloted our survey, how we found the participants, and how we analyzed the results. Section IV presents these results in detail, and Section V describes their specific implications. Finally, we describe the potential threats to the validity of our study in Section VI and conclude the paper in Section VII.
II Related Work
Recently, numerous research studies have investigated how introducing AI assistance for coding impacts programmers [14, 15, 16, 17, 18, 19, 11, 20, 21, 22, 23, 24, 25, 26]. This line of research aims to understand how programmers use, perceive, and benefit from AI assistants in different programming scenarios. In the following, we describe in more detail the works that constitute large-scale surveys aimed at understanding the advantages and challenges of AI usage for coding and delineate the distinctions between these surveys and our own research efforts.
@@ II Related Work
Amoozadeh et al. [22] discuss the exponential growth of AI systems and their potential for Computer Science Education, highlighting the influence of trust on student adoption and learning. They surveyed 253 students, eliciting feedback regarding trust, motivation, and confidence from AI users, as well as regarding interest and awareness from those who do not use AI. The authors found that trust was positively correlated with improving motivation, confidence, and knowledge in students regardless of their perception of confidence in their ability in programming.
Moreover, several companies from industry conducted large-scale studies to form an understanding of the market of AI tooling for coding. A survey by Stack Overflow [9] of 89,184 developers found that 70% are using or planning to use AI tools, with 77% expressing a favorable view. Respondents noted increased productivity with these tools, as well as their trust in their accuracy.
@@
In general, existing research on the utilization of AI assistants covers a range of perspectives, with some studies focusing on a single activity, such as code completion, while others offer more generalized insights that may overlook the nuances of different specific activities within software development. In contrast to this existing body of work, our study delves into developers’ perspectives regarding the use of AI assistance across distinct activities within the software development life-cycle and their individual stages, including a detailed analysis of their open-ended answers and concerns raised in them. Understanding developers’ perceptions within these specific activities is crucial as it enables us to identify existing challenges and concentrate our research efforts on addressing them effectively.
III Methodology
@@ 3. RQ3
III-A Survey Structure
Firstly, we ask the developers an array of general demographic questions (years of professional experience, employment status) and questions about their jobs (primary programming languages, job role, job level, and type of developed software). Next, to answer RQ1, we ask them whether they are familiar with specific AI assistant tools, given a list, and their overall opinion about the code generated by assistants.
@@ III-A Survey Structure
First, we ask the participants to rate the given activity on a 5-point Likert scale in three dimensions:
@@
Then, to answer RQ2, for each activity, we provide a list of stages, or component steps, within this activity and ask the participant whether they employ an AI assistant on this stage. Some stages are the same for all activities, for example, “Chatting with an AI assistant to brainstorm ideas”, while the majority are unique, like “Generating data and resources for tests (e.g., inputs and outputs)” for writing tests, or “Generating commit messages” for writing natural-language artifacts. The list of stages for each activity was compiled and polished during the pilot runs of the survey. Also, the participants could enter their own steps in the “Other” field.
To answer RQ3, for each activity, we ask an open-ended question about what prevents the participants from using AI assistants in the stages they did not select. Finally, we ask whether they would use an AI assistant in those steps if the issues mentioned in the previous question were resolved.
To ensure the robustness of our survey, we carried out five pilot runs with our colleagues: two software developers, two ML engineers, and one UX researcher, — and updated our survey based on their feedback. The full survey, including all the stages of activities, is available in our supplementary materials [13].
III-B Data Collection
The participants were gathered from both Industry and Academia communities. The survey link was advertised via personal social media by the authors of the paper and was emailed to the list of developers who are subscribed to the JetBrains products and who gave their consent for contact with research purposes. As a thank you, participants were given the opportunity to enter a draw for one of five 50 USD Amazon eGift Cards or an equivalent-value JetBrains product pack. The study was conducted in line with the company’s ethical standards, adhering to the values and guidelines outlined in the ICC/ESOMAR International Code [30].
III-C Data Analysis
After collecting the responses, we went on to analyze them. This includes filtering the results, calculating correlations, and processing open-ended questions.
Filtering results. At the end of the three weeks during which the survey remained open, we received 780 responses, of which 547 were complete.
@@ III-C Data Analysis
Processing open-ended questions (RQ3). For questions about why the respondents do not use AI assistants, we used thematic analysis [34] to group the responses. The first stage of the analysis was open-coding the responses. We decided to process all the responses together, going activity by activity: firstly, implementing new features, then writing tests, etc. This way, the obtained codes and later themes would be general enough, however, activity-specific codes would not be lost. The first two authors independently read all the responses and independently coded them. Then, the two authors discussed their codes until they reached a consensus for each response.
@@ IV-A Demographics
IV-B RQ1: General Usage and Perception of AI Assistants in SE
@@
IV-C RQ2: Activities and Their Stages
@@
TABLE I: Spearman correlations between answers to questions about Enjoyability, Easiness, and Delegability of each of the five main activities. The underlined values are statistically significant with $p<0.01$.
@@
Additionally, we explored the correlation between the difficulty of the activity and the willingness to delegate it. Overall, from Figure 2, it can be seen that opinions about the easiness of activities are more uniform, without as big differences as for other qualities. The results, presented in Table I support the idea that the correlation between the difficulty of the activity and willingness to delegate is weak and mostly not statistically significant, meaning that there is no connection between these two aspects of activity for our sample.
@@
Stages of Implementing new features. The first thing we can notice about Implementing new features is that only 12.7% of respondents—the lowest percentage among all activities—chose the option None of the above, meaning that they do not utilize AI tools for any of the stages in this activity. This means that the other 87.3% use AI in at least one of the stages of this activity, making it the most popular one for AI assistance.
@@
Stages of Writing natural-language artifacts. Similarly to Writing tests and Refactoring, here also about a quarter of respondents chose None of the above, and three-quarters of them use AI assistants in at least one stage of the activity.
@@
IV-D RQ3: Reasons for Not Using Assistants
@@
TABLE II: The resulting themes in the answers to the question “What prevents you from using AI assistants in the workflow steps you didn’t select above?”
* The percentage is calculated from all 1,559 coded responses across all activities.
@@
Theme Description Responses* Lack of need for AI assistance Respondents don’t try the assistants, express no need or no interest in them, state that the existing non-AI-based tools work well, etc. 22.5% AI-generated output is inaccurate The output generated by an AI assistant is inaccurate, the code is incorrect, the model hallucinates, etc. 17.7% User’s lack of trust and the desire to feel in control Respondents don’t trust the model, consider it to be unreliable, want to feel control over their code, believe that certain things have to be done by people, etc. 15.7% Lack of understanding context by AI assistant The assistant does not understand the context of the task or the underlying reason, cannot analyze the full codebase, cannot understand the requirements or business logic, etc. 14.4% User’s limited knowledge or understanding of AI assistants and their capabilities Respondents don’t know enough about the AI assistants or how they work, don’t know how to do the given task with it, don’t know what to ask, etc. 10.2% User’s desire to perform the task themselves and learn Respondents want to do the work themselves, love the process, want to understand the code better or to learn, etc. 7.3% Time inefficiency Using AI assistants does not save any time for the respondent, it takes longer to prompt the assistant or to later fix the output than to do the task themsevles, etc. 5.5% AI-generated output is not useful The output generated by an AI assistant is not useful, rudimentary, etc. 4.6% Company policies, NDAs, etc. Respondents are prohibited from using the AI assistants by their companies, NDAs, various policies, etc. 3.6% User’s negative attitude towards AI Respondents feel general negative emotions towards AI: hate, not wanting to rely too much, fear of being replaced, etc. 3.2% Lack of compliance with non-functional requirements in AI assistant’s output The output generated by an AI assistant does not fit the styling conventions of the project, has poor readability, is too long or too short, is not human-like, etc. 3.2% Legal and ethical considerations Respondents are worried about their privacy, about the copyright of the output, think that using AI is unethical, etc. 3.1% Challenges with AI integration and usability in development environments The AI assistant is not integrated well into the IDE, does not have the necessary functionality, has tooling and usability problems, etc. 3.0% Lack of access Respondents cannot access AI assistants, they are too expensive, etc. 2.6% Security concerns Respondents are worried about security, the AI assistants providing unsafe code, etc. 2.3% Limitations in AI’s creativity AI assistant lacks creativity, cannot provide complex solutions, provides generic code, etc. 2.2% Workflow disruption Respondents are not used to using AI assistants, have challenges with adopting it, feel like it breaks their workflow, etc. 2.1% Limitations of knowledge in the AI model The model in the AI assistant is outdated, lacks domain knowledge, is not trained well for non-English language, etc. 1.9% Inefficiency for specific tasks The AI assistant is incapable of carrying out some specific task, etc. 1.7% Challenges in communicating user intentions to AI assistant Respondents find it hard to communicate their intentions to the AI assistant, the assistant cannot follow instructions, requires directly pointing it to the necessary code, etc. 1.0%
@@
Finally, it is critical to understand why developers are hesitant to use AI assistants for some development activities. To do that, we studied their responses to open-ended questions about factors that prevent respondents from using AI assistants. Our thematic analysis resulted in 20 distinct themes. Their full list with detailed description is provided in Table II, with the percentage calculated from all 1,559 coded responses across all five activities. We first discuss the themes in general and then highlight more specific problems for each activity.
@@
In the following, we highlight specific issues mentioned in different activities. For each activity, we report the three themes, the relative percentage of which in that activity’s responses is the highest compared to the overall percentage presented in Table II, together with the reasoning that we collected from the responses.
@@
Another aspect important for implementing new features is Time inefficiency, raised in 11.3% of answers for this activity. Some respondents mentioned that correctly prompting the AI assistant takes longer than writing the code themselves, and some mentioned that fixing the generated code takes longer.
@@
Specifics of Writing tests. For Writing tests, as many as 32.1% responses mentioned Lack of need for AI assistance. Some respondents do not do testing at all, and some mentioned that the existing tools work well for finding untested areas of code. P98 wrote: “There are multiple extensions for both Rider and Visual Studio as well as commandline tools that show code coverage, so there’s no need to use AI for that.”
@@
Specifics of Bug triaging. 18.5% of responses related to Bug triaging mentioned Lack of understanding context by AI assistant. Respondents mentioned that finding and fixing non-trivial bugs requires a large context, both in terms of code and in terms of a more high-level, conceptual understanding of the program, which the AI assistants lack. P253 wrote: “A bug can come from a subpar execution of an idea or because a previously unknown behaviour of the library/programming language/api. Investigating and fixing such bugs requires understanding a great deal of information…”
@@
Specifics of Writing natural-language artifacts. 29% of responses about natural-language artifacts contained the theme of Lack of need for AI assistance. A lot of respondents did not try AI assistants for certain artifacts, especially larger ones like READMEs.

### v2

@@ Abstract
Context. The last several years saw the emergence of AI assistants for code — multi-purpose AI-based helpers in software engineering. As they become omnipresent in all aspects of software development, it becomes critical to understand their usage patterns.
Objective. We aim to better understand how specifically developers are using AI assistants, why they are not using them in certain parts of their development workflow, and what needs to be improved in the future.
Methods. In this work, we carried out a large-scale survey aimed at how AI assistants are used, focusing on specific software development activities and stages. We collected opinions of 481 programmers on five broad activities: (a) implementing new features, (b) writing tests, (c) bug triaging, (d) refactoring, and (e) writing natural-language artifacts, as well as their individual stages.
Results. Our results provide a novel comparison of different stages where AI assistants are used that is both comprehensive and detailed. It highlights specific activities that developers find less enjoyable and want to delegate to an AI assistant, e.g., writing tests and natural-language artifacts. We also determine more granular stages where AI assistants are used, such as generating tests and generating docstrings, as well as less studied parts of the workflow, such as generating test data. Among the reasons for not using assistants, there are general aspects like trust and company policies, as well as more concrete issues like the lack of project-size context, which can be the focus of the future research.
Conclusion. The provided analysis highlights stages of software development that developers want to delegate and that are already popular for using AI assistants, which can be a good focus for features aimed to help developers right now. The main reasons for not using AI assistants can serve as a guideline for future work.
keywords:
LLMs , AI Assistants , Software Development Lifecycle
journal: Information and Software Technology
1 Introduction
@@
Given their rapid growth, understanding software developers’ perceptions and needs regarding AI assistants is paramount in shaping the future of software development. Recent studies have started looking into why AI assistants are used, what their high-level drawbacks are [10, 11, 12], etc. However, the existing studies do not consider individual activities within the software development life-cycle and their specificities. Because of this, it is challenging to transform the results of such works into actionable implications, since they only deal with general concerns like quality and security, without taking into account their concrete manifestations in the developer’s workflow. Conversely, other works focus on individual narrow stages, e.g., code completion, test generation, etc., and study them in detail. While they are more specific, it is difficult to compare and prioritize their results, because they do not take the broader picture into account. The quick development of AI assistants requires a comprehensive yet granular analysis that considers the full software development lifecycle and its detailed steps at the same time, thus allowing to find areas where AI assistants are already employed to focus on them in the short term and identify shortcomings to address them in the future. To the best of our knowledge, no study has yet been conducted on such an analysis.
@@ 3. RQ3
  • Usage patterns. We find that 84.2% of the respondents occasionally or regularly use AI assistants, that Implementing new features is the most popular activity to use assistants, and that among individual stages, generating and summarizing code are the most widely used.
  • Areas of focus. Taking into account which activities the developers find less enjoyable and want to delegate, as well as what stages AI assistants are used on, we highlight areas where the research needs to focus on to bring value to users right now. This includes Writing tests in general and generating data and resources for tests in particular, as well as Writing natural language artifacts.
  • Reasons for not using AI. We provide 20 distinct themes that represent different reasons why developers are not using AI assistants and report the main ones for all activities. The most popular ones are Lack of need for AI assistance, AI-generated output being inaccurate, User’s lack of trust and the desire to feel in control, and Lack of understanding context by AI assistant.
  • Areas of future improvement. Based on the prevalence of different reasons, we formulate the main areas of future work that are needed to overcome the shortcomings that the respondents describe. This includes the improvement of base systems, better integration into developers’ workflow, and more actively educating and informing users about the assistants’ capabilities and limitations.
The detailed description of the survey, the collected responses, as well as some additional results that did not fully fit into the paper can be found in supplementary materials [13].
The remainder of this paper is organized as follows. In Section 2, we discuss the related work. In Section 3, we describe how we developed and piloted our survey, how we found the participants, and how we analyzed the results. Section 4 presents these results in detail, and Section 5 describes their specific implications. Finally, we describe the potential threats to the validity of our study in Section 6 and conclude the paper in Section 7.
2 Related Work
AI assistance is a very active area of research, and numerous studies have investigated its impact on how programmers code. In this section, we describe the main surveys and studies — both academic and industrial — that explore the users’ perceptions and benefits from AI assistants.
@@ 2 Related Work
Zhou et al. [11] aimed at understanding the issues and challenges developers face when using Copilot in practice, as well as their underlying causes and potential solutions. The results of the analysis of data from 476 GitHub issues, 706 GitHub discussions, and 142 Stack Overflow posts highlight that users commonly face Operation and Compatibility Issues with Copilot. The most frequent causes include Copilot Internal Issues, Network Connection Problems, and Editor/IDE Compatibility Issues. Predominant solutions involve Bug Fixes by Copilot, Configuration/Setting Modifications, and Using Suitable Versions. The authors suggest integrating a code explanation feature to enhance Copilot’s overall utility and effectiveness in practical development scenarios, considering the additional time required for code suggestion verification.
Pothukuchi et al. [16] examine the potential of generative AI to automate and enhance traditional software development practices, thereby improving efficiency and reducing costs. A total of 30 professionals participated in the interview study and highlighted significant improvements in development speed and code quality. The authors indicate that generative AI can significantly transform the SDLC by automating repetitive tasks and providing intelligent code suggestions, and propose a new model called Generative AI-Assisted Software Development Lifecycle.
Mozannar et al. [17] conducted a comprehensive study to inform the understanding of how developers interact with AI tools and how to improve this experience. They studied the impact of GitHub Copilot on programmers’ behavior during coding sessions. As a result, they identified 12 common programmer activities related to AI code completion systems. The researchers found that developers spend more time reviewing code than writing it.
Barke et al. [18] propose a grounded theory of the bimodal nature of programmers’ interactions with AI assistance that states that there are two main interaction modes between which developers fluidly switch while programming — exploration and acceleration. In acceleration mode, the programmer already knows what they want to do next, and AI helps them get there quicker; interactions in this mode are fast and do not break programmer’s flow. In exploration mode, the programmer is not sure how to proceed and uses Copilot to explore their options or get a starting point for the solution; interactions in this mode are slow and deliberate, and include explicit prompting and more extensive validation.
Moreover, several companies from industry conducted large-scale studies to form an understanding of the market of AI tooling for coding. A survey of 89,184 developers conducted in 2023 by Stack Overflow [9] found that 70% are using or planning to use AI tools, with 77% expressing a favorable view. Respondents noted increased productivity with these tools, as well as their trust in their accuracy. The most recent Stack Overflow survey [19], conducted on 1700 people, revealed the continuation of the trend — users report more quality work time.
@@
In general, existing research on the utilization of AI assistants covers a range of perspectives, with some studies focusing on a single activity, such as code completion, while others offer more generalized insights that may overlook the nuances of different specific activities within software development. In contrast to this existing body of work, our study aims to combine the strengths of both approaches to allow for a more direct comparison of various areas of application of AI assistants. Taking into account the entire software development lifecycle, while also asking about specific low-level work aspects, can allow our findings to be both concrete and comprehensive. Having a wide and detailed palette of opinions within one large-scale survey provides an opportunity to prioritize both short-term improvements to features that are already in use and long-term research into reasons why some features are not yet widely employed.
3 Methodology
@@ 3. RQ3
3.1 Survey Structure
The full survey, including all the stages of activities, is available in our supplementary materials [13].
Firstly, we ask the developers an array of general demographic questions (years of professional experience, employment status) and questions about their jobs (primary programming languages, job role, job level, and type of developed software). Next, to answer RQ1 and establish the general usage and perception of AI assistants, we inquire about the participants’ familiarity with specific popular AI assistant tools, as well as their opinions about the main qualities of the code generated by AI assistants: usability, accuracy, alignment with non-functional requirements, and security.
@@ 3.1 Survey Structure
First, to answer RQ2 and discover how participants perceive the activities, we ask them to rate each activity on a 5-point Likert scale in three dimensions:
@@
Then, we deepen our understanding of activities by considering their individual stages, which allows us to make the results of this research not only comprehensive but also detailed. For each activity, we provide a list of stages, or component steps, within this activity and ask the participant whether they employ an AI assistant on this stage. Some stages are the same for all activities, for example, “Chatting with an AI assistant to brainstorm ideas”, while the majority are unique, like “Generating data and resources for tests (e.g., inputs and outputs)” for writing tests, or “Generating commit messages” for writing natural-language artifacts. Also, the participants could enter their own steps in the “Other” field.
Finally, to answer RQ3, for each activity, we ask an open-ended question about what prevents the participants from using AI assistants in the stages they did not select and whether they would use an AI assistant in those steps if the issues were resolved. Responses to these questions can allow us to not only quantitatively determine the least popular stages, but also establish distinct reasons for not using assistants there, thus leading the way to fixing them.
To ensure the robustness of our survey, we consulted experts in survey design and carried out five pilot runs with our colleagues. The experts included members of the Surveys Team and the Research Core team at JetBrains, who specialize in designing surveys and other forms of quantitative and qualitative research. The list of pilot participants was also diverse — two software developers, two ML engineers, and one UX researcher. Overall, the feedback from the experts and the pilot participants was used to update and refine the survey. In particular, we compiled and polished the list of stages for each activity. Also, the participants highlighted the need to have “Chatting with an AI assistant to brainstorm” as an option for each activity, as well as having an explicit “None” option.
3.2 Data Collection
The participants were invited from both Industry and Academia communities. The survey link was advertised via personal social media by the authors of the paper, three of whom are the employees of JetBrains Research, and was emailed to the list of people curated by JetBrains who gave their consent for contact with research purposes. Importantly, not all people in this mailing list are customers of JetBrains products, and the list includes the users of various IDEs and AI assistants. As a thank you, participants were given the opportunity to enter a draw for one of five 50 USD Amazon eGift Cards or an equivalent-value JetBrains product pack. The study was conducted in line with the company’s ethical standards, adhering to the values and guidelines outlined in the ICC/ESOMAR International Code [23].
3.3 Data Analysis
After collecting the responses, we analyzed them. This includes filtering the results, calculating correlations, and processing open-ended questions.
Filtering results. At the end of the three weeks during which the survey remained open, we received 780 responses, of which 547 were complete. Among these 547 responses, 542 came from the mailing list, indicating a strong industrial backing of our sample. This allowed us to collect diverse opinions from experienced developers acquainted with a variety of different tools, as evidenced further in Sections 4.1 and 4.2.
@@ 3.3 Data Analysis
Processing open-ended questions (RQ3). For questions about why the respondents do not use AI assistants, we used thematic analysis [27] to group the responses. The first stage of the analysis was open-coding the responses. We decided to process all the responses together, going activity by activity: firstly, implementing new features, then writing tests, etc. This way, the obtained codes and later themes would be general enough, however, activity-specific codes would not be lost. The first two authors independently read all the responses and independently coded them. Then, the two authors discussed their codes until they reached a consensus for each response using a negotiated agreement approach [28].
@@ 4.1 Demographics
4.2 RQ1: General Usage and Perception of AI Assistants in SE
@@
4.3 RQ2: Activities and Their Stages
@@
Table 1: Spearman correlations between answers to questions about Enjoyability, Easiness, and Delegability of each of the five main activities. The underlined values are statistically significant with $p<0.01$.
@@
Additionally, we explored the correlation between the difficulty of the activity and the willingness to delegate it. Overall, from Figure 2, it can be seen that opinions about the easiness of activities are more uniform, without as big differences as for other qualities. The results, presented in Table 1 support the idea that the correlation between the difficulty of the activity and willingness to delegate is weak and mostly not statistically significant, meaning that there is no connection between these two aspects of activity for our sample.
@@
Stages of Implementing new features. The first thing we can notice about Implementing new features is that only 12.7% of respondents — the lowest percentage among all activities — chose the option None of the above, meaning that they do not utilize AI tools for any of the stages in this activity. This means that the other 87.3% use AI in at least one of the stages of this activity, making it the most popular one for AI assistance. Interestingly, in Figure 3 we have seen that the smallest percentage of developers wanted to delegate Implementing new features, yet in reality the largest percentage of them use AI Assistants for some of the aspects of this activity. This can be explained by the fact that developers assume full delegation that they do not want, but actually employ assistants for some stages.
@@
Stages of Writing natural-language artifacts. Similarly to Writing tests and Refactoring, here also about a quarter of respondents chose None of the above, and three-quarters of them use AI assistants in at least one stage.
@@
4.4 RQ3: Reasons for Not Using Assistants
@@
Table 2: The resulting themes in the answers to the question “What prevents you from using AI assistants in the workflow steps you didn’t select above?”. * The percentage is calculated from all 1,559 coded responses across all activities.
@@
Theme Description Responses* Lack of need for AI assistance Respondents do not try the assistants, express no need or no interest in them, state that the existing non-AI-based tools work well, etc. 22.5% AI-generated output is inaccurate The output generated by an AI assistant is inaccurate, the code is incorrect, the model hallucinates, etc. 17.7% User’s lack of trust and the desire to feel in control Respondents do not trust the model, consider it to be unreliable, want to feel control over their code, believe that certain things have to be done by people, etc. 15.7% Lack of understanding context by AI assistant The assistant does not understand the context of the task or the underlying reason, cannot analyze the full codebase, cannot understand the requirements or business logic, etc. 14.4% User’s limited knowledge or understanding of AI assistants and their capabilities Respondents do not know enough about the AI assistants or how they work, do not know how to do the given task with it, do not know what to ask, etc. 10.2% User’s desire to perform the task themselves and learn Respondents want to do the work themselves, love the process, want to understand the code better or to learn, etc. 7.3% Time inefficiency Using AI assistants does not save any time for the respondent, it takes longer to prompt the assistant or to later fix the output than to do the task themsevles, etc. 5.5% AI-generated output is not useful The output generated by an AI assistant is not useful, rudimentary, etc. 4.6% Company policies, NDAs, etc. Respondents are prohibited from using the AI assistants by their companies, NDAs, various policies, etc. 3.6% User’s negative attitude towards AI Respondents feel general negative emotions towards AI: hate, not wanting to rely too much, fear of being replaced, etc. 3.2% Lack of compliance with non-functional requirements in AI assistant’s output The output generated by an AI assistant does not fit the styling conventions of the project, has poor readability, is too long or too short, is not human-like, etc. 3.2% Legal and ethical considerations Respondents are worried about their privacy, about the copyright of the output, think that using AI is unethical, etc. 3.1% Challenges with AI integration and usability in development environments The AI assistant is not integrated well into the IDE, does not have the necessary functionality, has tooling and usability problems, etc. 3.0% Lack of access Respondents cannot access AI assistants, they are too expensive, etc. 2.6% Security concerns Respondents are worried about security, the AI assistants providing unsafe code, etc. 2.3% Limitations in AI’s creativity AI assistant lacks creativity, cannot provide complex solutions, provides generic code, etc. 2.2% Workflow disruption Respondents are not used to using AI assistants, have challenges with adopting it, feel like it breaks their workflow, etc. 2.1% Limitations of knowledge in the AI model The model in the AI assistant is outdated, lacks domain knowledge, is not trained well for non-English language, etc. 1.9% Inefficiency for specific tasks The AI assistant is incapable of carrying out some specific task, etc. 1.7% Challenges in communicating user intentions to AI assistant Respondents find it hard to communicate their intentions to the AI assistant, the assistant cannot follow instructions, requires directly pointing it to the necessary code, etc. 1.0%
@@
Finally, it is critical to understand why developers are hesitant to use AI assistants for some development activities. To do that, we studied their responses to open-ended questions about factors that prevent respondents from using AI assistants. Our thematic analysis resulted in 20 distinct themes. Their full list with detailed description is provided in Table 2, with the percentage calculated from all 1,559 coded responses across all five activities. We first discuss the themes in general and then highlight more specific problems for each activity.
@@
In the following, we highlight specific issues mentioned in different activities. For each activity, we report the three themes, the relative percentage of which in that activity’s responses is the highest compared to the overall percentage presented in Table 2, together with the reasoning that we collected from the responses.
@@
Another aspect important for new features is Time inefficiency, raised in 11.3% of answers for this activity. Some respondents mentioned that correctly prompting the AI assistant takes longer than writing the code themselves, and some mentioned that fixing the generated code takes longer.
@@
Specifics of Writing tests. For Writing tests, as many as 32.1% responses mentioned Lack of need for AI assistance. Some respondents do not do testing at all, and some mentioned that the existing tools work well for finding untested areas of code. P98 wrote: “There are multiple extensions for both Rider and Visual Studio as well as commandline tools that show code coverage, so there is no need to use AI for that.”
@@
Specifics of Bug triaging. 18.5% of responses related to Bug triaging mentioned Lack of understanding context by AI assistant. Respondents mentioned that finding and fixing non-trivial bugs requires a large context, both in terms of code and in terms of a more high-level, conceptual understanding of the program, which the AI assistants lack. P253 wrote: “A bug can come from a subpar execution of an idea or because a previously unknown behavior of the library/programming language/api. Investigating and fixing such bugs requires understanding a great deal of information…”
@@
Specifics of Writing natural-language artifacts. 29% of responses about natural-language artifacts contained the theme of Lack of need for AI assistance. Interestingly, Section 4.3 and Figure 3 indicated that developers strongly want to delegate this activity. However, a lot of respondents said that they did not try AI assistants for certain artifacts, especially larger ones like READMEs. Moreover, some of them indicated that the existing tools work fine for them.